Two firms are producing identical goods in a market characterized by the inverse demand curve P = 120 – 4Q, where Q is the sum of Firm 1's and Firm 2's output, q1 + q2. Each firm's marginal cost is constant at $20.

Graph the reaction function for each firm and indicate the Nash equilibrium.
Answer #1

P= 120-4G and me, = mc2 = 20 We can determine the reaction function for firm 1 as follows. To maximize profit, it sets mRi= mSetting mRi= me, we get 120-891 - 492 = 20 → 89, = 100 - 492 191 = 12.5-0.592 firm 1s reaction I curve The same appling caku